# MACTAS: Self-Attention-Based Module for Inter-Agent Communication in Multi-Agent Reinforcement Learning

### v1

@@ 3.4 Exploration
Table 1: Basic experimental settings and hyperparameters. 1 for 2.5 mln. step maps. 2 for 7.5 mln. step maps.
@@
Table 2: MACTAS shape for each scenario.
Figure 3: Percentage of wins in test games in training time for the connection protocols MACTAS and MAIC, and the $Q$ architectures QMIX, QPLEX, and VDN.
Figure 4: The impact of our exploration scheme on MACTAS+QMIX.
Figure 5: The impact of our residual connection on MACTAS+QMIX.
4 Experiments
In this section, we present the results of our proposed MACTAS communication protocol in the StarCraft Multi-Agent Challenge (SMAC: Samvelyan et al. 2019) environment. We compare it with the state-of-the-art MAIC (Yuan et al. 2022), in combinations with 3 mixer algorithms: QMIX, QPLEX, VDN, and either MAIC or MACTAS communication protocol, on 6 maps. We present results of two ablation studies regarding the design of neural network connectivity and the exploration form. Additionally, in the Appendix, we report the sensitivity study of MACTAS to hyperparameters, lower-level SMAC settings, and analyses of the most interesting strategies that have emerged in training.
4.1 StarCraft Multi-Agent Challenge
StarCraft Multi-Agent Challenge (SMAC) is a micromanagement benchmark where two opposite teams face each other in different scenarios (maps), one team being controlled by the RL agent and the other by a built-in algorithm. Maps differ by difficulty, mainly due to an imbalance in the number of controlled units. For our experiments, we selected 2 hard and 4 super-hard scenarios, which are widely benchmarked in the MARL community:
@@ 4.1 StarCraft Multi-Agent Challenge
  • MMM2 scenario with a diverse units containing 1 medivac (medic), 2 marauders, and 7 marines fighting team with 1 more marauder and 1 more marine. Level: super-hard.
@@
In order to verify versatility of MACTAS algorithm, we run experiments on a diverse set of hard and super-hard scenarios. Every map provides the opportunity to find a unique winning strategy. The agent requires learning basic skills such as navigating the battlefield, managing health and shield, or focusing the fire. All considered scenarios are imbalanced in favor of the enemies; therefore, simple strategies are not sufficient to win the episode. Popular observed winning policies include crossing the obstacle on 2c_vs_64zg, flanking on 27m_vs_30m, ’kiting’ on 6h_vs_8z, ’hit and run’ on 3s5z_vs_3s6z (Samvelyan et al. 2019).
4.2 Experimental setting
We borrow the basic settings from NA2Q (Liu, Zhu, and Chen 2023) and extend their code with the implementation of MACTAS and MAIC. The algorithms are trained using the Double Q-Learning (Hasselt, Guez, and Silver 2016) algorithm. Each experiment is run for a fixed number of training episode steps – 2.5 million for 2c_vs_64zg, 5m_vs_6m, MMM2, 27m_vs_30m environments and 7.5 million for 6h_vs_8z, 3s5z_vs_3s6z environments, interleaved with test episodes every 2000 steps. Each experiment is repeated for seeds $1,\dots,5$. We report the average percentage of wins (and standard deviations in the Appendix) in test games.
@@ 4.2 Experimental setting
For MAIC communication, we use the same hyperparameters as in the original paper.
@@ 4.3 Results
On the 2c_vs_64zg map, all algorithms except MAIC+VDN achieve the same final performance. In terms of the area under the learning curve MACTAS performs similarly to MAIC for the QMIX and VDN mixers and better for the QPLEX mixer. MACTAS+QMIX and MAIC+QMIX give jointly the best results on this map. On the 5m_vs_6m map MACTAS+QPLEX achieves the same final result as MAIC+QPLEX but with a better area under the learning curve. MACTAS+QMIX gives the same final result as MAIC+QMIX, but has a worse area under the learning curve (MAIC+QMIX drops around two million steps). MACTAS+VDN achieves a better final result than MAIC+VDN with a similar area under the learning curve (MAIC+VDN starts to learn faster but is unable to go beyond $60\%$ of the mean of the won battles). On this map MACTAS+QPLEX gives the best result. We observe that on the 27m_vs_30m map, our algorithm achieves a better area under the learning curve for all three mixers, and for QPLEX, it also achieves a better final performance. For the other mixers, the final performance is similar. We hypothesize that the good performance of our algorithm is the case because the number of parameters of our algorithm is independent of the number of agents, and thus our method scales well in environments with a large number of agents. Overall, MACTAS+QMIX gives the best result on this map. On the MMM2 map, MAIC performs better for all three mixers in terms of the final result. However, the areas under the learning curves are similar and for MACTAS+QMIX and for MAIC+QMIX, as our algorithm starts to learn faster. MAIC+QPLEX performs the best on this map. This is the only map on which a version of our algorithm does not give the best result. On the 6h_vs_8z map MACTAS performs better than MAIC for the QMIX mixer in terms of the area under the learning curve and the final result. MACTAS+VDN gives a similar final result as MAIC+VDN, but with a better area under the learning curve. MACTAS+QPLEX achieves a similar final result as MAIC+QPLEX, but with a worse area under the learning curve. Overall, MACTAS+QMIX and MAIC+QPLEX give jointly the best result on this map. On the 3s5z_vs_3s6z map, MACTAS+QMIX and MACTAS+VDN give better results in terms of both the area under the learning curve and the final result than MAIC+QMIX and MAIC+VDN, respectively. MACTAS+QPLEX gives a similar final result to MAIC+QPLEX, but with a worse area under the learning curve. MACTAS+VDN performs the best on this map.
4.4 Ablations
@@
Secondly, we verify the role of our proposed exploration scheme. Figure 4 shows the impact of our exploration scheme for MACTAS+QMIX. We observe that our exploration method improves the results on the 3s5z_vs_3s6z map and is indifferent on the 27m_vs_30m and MMM2 maps. We used softmax temperature $0.33$ for the 3s5z_vs_3s6z map and $0.25$ for the 27m_vs_30m and MMM2 maps.
4.5 Discussion
@@
A moderate sensitivity of the performance of our proposed communication module to its shape requires searching for its best setting in additional runs. This slows the deployment of MACTAS.
5 Conclusions and Future Work
@@
We observe that $\epsilon$-greedy exploration applied independently by agents in MARL leaves room for improvement since it condones the agents’ policies falling into suboptimal Nash equilibria. We addressed this issue by proposing a combination of $\epsilon$-greedy and Boltzmann exploration, which increased the performance of MACTAS+QMIX on the 3s5z_vs_3s6z SMAC map. However, we believe that this issue requires further research and intend to undertake it in the future.
Acknowledgements
@@
Appendix A StarCraft Multi-Agent Challenge (SMAC)
As a benchmark for MACTAS, we chose the StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al. 2019), which is a simplified version of the StarCraft II game. The official repository is available on https://github.com/oxwhirl/smac. SMAC uses PySc2 to communicate with the game engine https://github.com/google-deepmind/pysc2. SMAC is characterized by decentralized management of each unit, and each unit is an independent agent. This feature allows for the testing of decentralized algorithms, including those originating from MARL. Benchmark scenarios are limited to combat with a limited number of units. Those scenarios do not account for the economic part of the game, the so-called “macro” part of the game – collecting resources and building structures. Resources in the game (gas and minerals) are required to build units and structures. Structures allow for the recruitment of units and their upgrading. The considered scenarios do not involve “worker” units, which collect resources. Instead, the scenarios focus on ’micro,’ the term used in the StarCraft II community for unit control during combat. The challenge for a human player is multitasking between these two tasks – developing the game economy and controlling the fighting units. SMAC focuses only on combat, thereby ensuring this benchmark repeatability, as each iteration involves a fight between the same specification units, eliminating the uncertainty with unit upgrades and different strategies for the economic part of the game. In StarCraft II, there are three races, and our experiments covered controlling units of all of them. The way to run SMAC is described in our repository.
@@ Appendix B Detailed Results
Figures 6, 7, and 8 contain more detailed results of our experiments described in Section 4.3. The curves on the plots represent the mean percentage test game win rate during the algorithm training for 6 different environments and various mixers. The shaded area around the curves is the standard deviation of the results among 5 different seeds (1 to 5) of the random number generator. For each seed, the mean percentage test win rate was calculated based on 32 independent test episodes. We observe robustness of MACTAS with the QMIX mixer – on the most challenging environments 6h_vs_8z and 3s5z_vs_3s6z, it outperforms MAIC+QMIX by a large margin; on 27m_vs_30m, it learns much faster; on 2c_vs_64zg and 3s5z_vs_3s6z, it performs similarly, and only slightly worse on the MMM2 map. For the QPLEX mixer, our method performs much better than MAIC on the 2c_vs_64zg and 27m_vs_30m maps, slightly better on the 5m_vs_6m, slightly worse on MMM2, learns slower but achieves similar final result on the 6h_vs_8z and 3s5z_vs_3s6z maps. Our method combined with the VDN mixer gives significantly better results than MAIC on the 3s5z_vs_3s6z map, slightly better on 6h_vs_8z, 27m_vs_30m, 5m_vs_6m, marginally better on the 2c_vs_64zg map and marginally worse on the MMM2 map.
Figure 6: Means and standard deviations of the percentage wins in the test games during training for QMIX with MAIC, MACTAS, and bare.
Figure 7: Means and standard deviations of the percentage wins in the test games during training for QPLEX with MAIC, MACTAS, and bare.
Figure 8: Means and standard deviations of the percentage wins in the test games during training for VDN with MAIC, MACTAS, and bare.
@@ Appendix C Tactics used by MACTAS
The SMAC challenge forces agents to learn various tactics, such as attacking from distance melee units (’kiting’) or using movement advantage (’hit and run’). In this section, we provide an overview of the tactics used by our model based on our understanding of StarCraft II. We observed differences in the behavior of our model depending on the random seed. The description of tactics is based on replays of our models for seed = 1.
C.1 2c_vs_64zg
Figure 9: 2c_vs_64zg replay.
@@
C.2 27m_vs_30m and 5m_vs_6m
Figure 10: 27m_vs_30m replay.
Figure 11: 5m_vs_6m replay.
The 27m_vs_30m and the 5m_vs_6m scenarios consist of marine units only. They are similar in the way that the opposing force numerically outnumbers the agents under control. In Figure 10, the agent’s units place themselves in a semicircle formation, also known as flanking, which maximizes the number of shooting units, while the enemies block each other. Units controlled by MACTAS focus fire on enemy units with lower health, in order to quickly reduce the number of opponents. Furthermore, units with low health retreat and continue fighting, while avoiding being shot. This behavior is presented in Figure 11: Agents with low health retreat when others focus fire on one enemy unit. The winning strategy is similar for both 5m_vs_6m and 27m_vs_30m. 5m_vs_6m is labeled as a hard benchmark, while 27m_vs_30m is classified as super-hard. The percentage imbalance between units is higher in 5m_vs_6m. However, 27m_vs_30m has a larger number of units that need to cooperate.
C.3 MMM2
Figure 12: MMM2 replay.
The MMM2 is a scenario with a diverse set of units containing 1 medivac (medic), 2 marauders, and 7 marines fighting a team with 1 more marauder and 1 more marine. Medivac cannot attack, but can heal units as long as it has energy. MACTAS ground units focus fire on enemy units and eliminate them one by one, and fall back when their health is low. The enemy medivac does not change healed unit, while the MACTAS’ medivac changes healed unit to the one with a critically low health or to the currently attacked unit. Units trained with MACTAS move medivac to the front line and focus enemy fire on it, and when its health level is low, retreat it. MACTAS’ medivac taking enemy fire secures outnumbered marines and marauders. Typically, human players try to save their medivacs because they are used for healing units between fights during a regular game. Only under certain conditions do they sacrifice their medivacs; for example, when they are breaking through an opponent’s line of defense to focus enemy fire on the medivacs and away from damage-dealing units. In this scenario, we can observe similar behavior in the MACTAS medivac control, which results in winning combat against an opponent who has a numerical advantage.
C.4 6h_vs_8z
Figure 13: 6h_vs_8z replay.

### v2

@@ 3.4. Exploration
Table 1. Basic experimental settings and hyperparameters. 1 for 2.5 mln. step maps. 2 for 7.5 mln. step maps.
@@
All SMACv2 scenarios 3 512
Table 2. MACTAS shape for each scenario.
Figure 3. Percentage of wins in test games in training time for the connection protocols MACTAS and MAIC, and the $Q$ architectures QMIX, QPLEX, and VDN.
Figure 4. Percentage of wins and standard deviations in test games in training time for the connection protocols MACTAS, MAIC, and the bare mixer for the $Q$ architecture QMIX
Figure 5. The impact of our exploration scheme on MACTAS+QMIX.
Figure 6. The impact of our residual connection on MACTAS+QMIX.
4. Experiments
In this section, we present the results of our proposed MACTAS communication protocol in the StarCraft Multi-Agent Challenge environments (SMAC Samvelyan et al. (2019) and SMACv2 Ellis et al. (2023)). We compare it with the state-of-the-art MAIC Yuan et al. (2022), in combinations with 3 mixer algorithms: QMIX, QPLEX, and VDN, on 6 maps. We present results of two ablation studies regarding the design of neural network connectivity and the exploration form. Additionally, in the Appendix, we report the sensitivity study of MACTAS to hyperparameters, lower-level SMAC and SMACv2 settings, and analyses of the most interesting strategies that have emerged in training.
4.1. StarCraft Multi-Agent Challenge
The StarCraft Multi-Agent Challenge (SMAC and SMACv2) is a micromanagement benchmark where two opposing teams face each other in different scenarios (maps), with one team controlled by the RL agent and the other by a built-in algorithm. Maps differ by difficulty, mainly due to an imbalance in the number of controlled units. For our experiments, we selected 2 hard and 4 super-hard SMAC scenarios, which are widely benchmarked in the MARL community:
@@ 4.1. StarCraft Multi-Agent Challenge
  • MMM2 scenario with diverse units containing 1 medivac (medic), 2 marauders, and 7 marines fighting team with 1 more marauder and 1 more marine. Level: super-hard.
@@
To verify the versatility of the MACTAS algorithm, we run experiments on a diverse set of hard and super-hard scenarios. Every map provides the opportunity to find a unique winning strategy. The agent requires learning basic skills, such as navigating the battlefield, managing health and shield, or focusing fire. All considered scenarios are imbalanced in favor of the enemies; therefore, simple strategies are insufficient to win the episode. Popular observed winning policies include crossing the obstacle on 2c_vs_64zg, flanking on 27m_vs_30m, ’kiting’ on 6h_vs_8z, ’hit and run’ on 3s5z_vs_3s6z Samvelyan et al. (2019).
In addition to classical SMAC, we benchmark the algorithms in its newer iteration, SMACv2 Ellis et al. (2023), which introduces new maps focused on scenario randomization. Each of the new maps defines the number of units in each team and the race (the same for both teams), while the exact types and starting positions are randomized according to the configuration. We use the default configuration for each scenario recommended in the SMACv2 repository. Our experiments focus on more challenging configurations, where the player team is outnumbered by the enemy team controlled by in-game AI, with 6 tested scenarios: zerg_5_vs_5, zerg_10_vs_11, zerg_20_vs_23, protoss_5_vs_5, protoss_10_vs_11, and protoss_20_vs_23.
4.2. Experimental setting
We borrow the basic settings from NA2Q Liu et al. (2023b) and extend their code with the implementation of MACTAS and MAIC. The algorithms are trained using the Double Q-Learning Hasselt et al. (2016) algorithm. Each experiment is run for a fixed number of training episode steps – 2.5 million for 2c_vs_64zg, 5m_vs_6m, MMM2, 27m_vs_30m environments and 7.5 million for 6h_vs_8z, 3s5z_vs_3s6z and all SMACv2 environments, interleaved with test episodes every 2000 steps. Each experiment is repeated for seeds $1,\dots,5$. We report the average percentage of wins in test games, along with the standard deviations in the Appendix.
@@ 4.2. Experimental setting
For MAIC communication, we employ the same hyperparameters as those used in the original paper.
@@ 4.3. Results
On the 2c_vs_64zg map, all algorithms except MAIC+VDN achieve the same final performance. In terms of the area under the learning curve MACTAS performs similarly to MAIC for the QMIX and VDN mixers and better for the QPLEX mixer. MACTAS+QMIX and MAIC+QMIX give jointly the best results on this map. On the 5m_vs_6m map MACTAS+QPLEX achieves the same final result as MAIC+QPLEX but with a better area under the learning curve. MACTAS+QMIX gives the same final result as MAIC+QMIX, but has a worse area under the learning curve (MAIC+QMIX drops around two million steps). MACTAS+VDN achieves a better final result than MAIC+VDN with a similar area under the learning curve (MAIC+VDN starts to learn faster but is unable to go beyond $60\%$ of the mean of the won battles). On this map MACTAS+QPLEX gives the best result. We observe that on the 27m_vs_30m map, our algorithm achieves a better area under the learning curve for all three mixers, and for QPLEX, it also achieves a better final performance. For the other mixers, the final performance is similar. We hypothesize that the good performance of our algorithm is because the number of parameters in our algorithm is independent of the number of agents. Thus, our method scales well in environments with a large number of agents. Overall, MACTAS+QMIX gives the best result on this map. On the MMM2 map, MAIC performs better for all three mixers in terms of the final result. However, the areas under the learning curves are similar and for MACTAS+QMIX and for MAIC+QMIX, as our algorithm starts to learn faster. MAIC+QPLEX performs the best on this map. This is the only map on which a version of our algorithm does not give the best result. On the 6h_vs_8z map MACTAS performs better than MAIC for the QMIX mixer in terms of the area under the learning curve and the final result. MACTAS+VDN gives a similar final result as MAIC+VDN, but with a better area under the learning curve. MACTAS+QPLEX achieves a similar final result as MAIC+QPLEX, but with a worse area under the learning curve. Overall, MACTAS+QMIX and MAIC+QPLEX give jointly the best result on this map. On the 3s5z_vs_3s6z map, MACTAS+QMIX and MACTAS+VDN give better results in terms of both the area under the learning curve and the final result than MAIC+QMIX and MAIC+VDN, respectively. MACTAS+QPLEX gives a similar final result to MAIC+QPLEX, but with a worse area under the learning curve. MACTAS+VDN performs the best on this map.
SMACv2
The basic results are presented in Fig. 4. We compare MACTAS with MAIC and the bare mixer on 6 maps and the QMIX mixer. In all scenarios, our algorithm outperforms MAIC and the bare mixer. We observe that the differences between MACTAS+QMIX and MAIC+QMIX are similar to the differences between MAIC+QMIX and the bare mixer.
4.4. Ablations
@@
Secondly, we verify the role of our proposed exploration scheme. Figure 5 shows the impact of our exploration scheme for MACTAS+QMIX. We observe that our exploration method improves the results on the most challenging SMAC maps, which require $7.5$ million training steps: 3s5z_vs_3s6z, and 6h_8z. It is also indifferent on the maps: 2c_vs_64zg, 27m_vs_30m, and MMM2. Our exploration scheme slows down the training for the 5m_vs_6m map. We used softmax temperature $0.33$ for the map 3s5z_vs_3s6z and $0.25$ for the maps: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, and 6h_8z.
4.5. Discussion
@@
A moderate sensitivity of the performance of our proposed communication module to its shape requires searching for its best setting in additional runs. This slightly slows down the deployment of MACTAS.
5. Conclusions and Future Work
@@
We observe that $\epsilon$-greedy exploration applied independently by agents in MARL leaves room for improvement since it condones the agents’ policies falling into suboptimal Nash equilibria. We addressed this issue by proposing a combination of $\epsilon$-greedy and Boltzmann exploration, which increased the performance of MACTAS+QMIX on the 6h_vs_8z, and the 3s5z_vs_3s6z SMAC maps. However, we believe that this issue requires further research and intend to undertake it in the future.
{acks}
@@
Appendix A StarCraft Multi-Agent Challenge (SMAC & SMACv2)
As a benchmark for MACTAS, we chose the StarCraft Multi-Agent Challenge (SMAC) Samvelyan et al. (2019) and its newer version SMACv2 Ellis et al. (2023), which are simplified versions of the StarCraft II game. The framework exists in two versions – SMAC (code available at https://github.com/oxwhirl/smac), which introduced predefined deterministic scenarios that often include purposeful selection of units or obstacles placement, and SMACv2 (code available at https://github.com/oxwhirl/smacv2), which added new randomized scenarios based on plain maps and randomized units generation. From here onward, we will use ”SMACv1” to refer to the first version of the benchmark and ”SMAC” to refer to both SMACv1 and SMACv2, making an explicit distinction where necessary. SMAC uses PySc2 to communicate with the game engine https://github.com/google-deepmind/pysc2. SMAC is characterized by the decentralized management of each unit, where each unit acts as an independent agent. This feature allows for the testing of decentralized algorithms, including those originating from MARL. Benchmark scenarios are limited to combat with a limited number of units. Those scenarios do not account for the economic aspect of the game, the so-called “macro” part of the game – collecting resources and building structures. Resources in the game (gas and minerals) are required to build units and structures. Structures enable the recruitment of units and their subsequent upgrading. The considered scenarios do not involve “worker” units, which collect resources. Instead, the scenarios focus on ’micro’, a term used in the StarCraft II community to describe unit control during combat. In SMACv2, the new addition is randomized maps defined by race (same for both teams), number of units in each team, thus map names like zerg_10_vs_11. The new maps randomize types of units according to the probability of each unit type defined in scenario configuration files called capabilities config. In addition, the configuration allows for specifying the units’ start position, their spread, and configuration – either surround, reflect, or random probability of each. The challenge for a human player is to multitask between these two tasks – developing the game economy and controlling the fighting units. SMAC focuses solely on combat, thereby ensuring repeatability of this benchmark, as each iteration involves a fight between the same specification units, eliminating uncertainty due to unit upgrades and different strategies in the economic part of the game. In StarCraft II, there are three races, and our experiments covered controlling units of all of them. The process for running SMAC is described in our repository.
@@ Appendix B Detailed Results
Figures 7, 8, and 9 contain more detailed results of our experiments described in Section 4.3. The curves on the plots represent the mean percentage test game win rate during the algorithm training for 6 different environments and various mixers. The shaded area around the curves is the standard deviation of the results among 5 different seeds (1 to 5) of the random number generator. For each seed, the mean percentage test win rate was calculated based on 32 independent test episodes. We observe robustness of MACTAS with the QMIX mixer – on the most challenging environments 6h_vs_8z and 3s5z_vs_3s6z, it outperforms MAIC+QMIX by a large margin; on 27m_vs_30m, it learns much faster; on 2c_vs_64zg and 3s5z_vs_3s6z, it performs similarly, and only slightly worse on the MMM2 map. For the QPLEX mixer, our method performs much better than MAIC on the 2c_vs_64zg and 27m_vs_30m maps, slightly better on the 5m_vs_6m, slightly worse on MMM2, learns slower but achieves similar final result on the 6h_vs_8z and 3s5z_vs_3s6z maps. Our method combined with the VDN mixer gives significantly better results than MAIC on the 3s5z_vs_3s6z map, slightly better on 6h_vs_8z, 27m_vs_30m, 5m_vs_6m, marginally better on the 2c_vs_64zg map, and marginally worse on the MMM2 map.
Figure 7. Means and standard deviations of the percentage wins in the test games during training for QMIX with MAIC, MACTAS, and bare.
Figure 8. Means and standard deviations of the percentage wins in the test games during training for QPLEX with MAIC, MACTAS, and bare.
Figure 9. Means and standard deviations of the percentage wins in the test games during training for VDN with MAIC, MACTAS, and bare.
@@ Appendix C Tactics used by MACTAS
The SMAC challenge forces agents to learn various tactics, such as attacking from a distance with melee units (’kiting’) or using a movement advantage (’hit and run’). In this section, we provide an overview of the tactics used by our model based on our understanding of StarCraft II. We observed differences in the behavior of our model depending on the random seed. The description of tactics is based on replays of our models for a seed value of 1.
C.1. 2c_vs_64zg
Figure 10. 2c_vs_64zg replay.
@@
C.2. 27m_vs_30m and 5m_vs_6m
Figure 11. 27m_vs_30m replay.
Figure 12. 5m_vs_6m replay.
The 27m_vs_30m and the 5m_vs_6m scenarios consist of marine units only. They are similar in the way that the opposing force numerically outnumbers the agents under control. In Figure 11, the agent’s units place themselves in a semicircle formation, also known as flanking, which maximizes the number of shooting units, while the enemies block each other. Units controlled by MACTAS focus fire on enemy units with lower health to quickly reduce the number of opponents. Furthermore, units with low health retreat and continue fighting, while avoiding being shot. This behavior is presented in Figure 12: Agents with low health retreat when others focus fire on one enemy unit. The winning strategy is similar for both 5m_vs_6m and 27m_vs_30m. 5m_vs_6m is labeled as a hard benchmark, while 27m_vs_30m is classified as super-hard. The percentage imbalance between units is higher in 5m_vs_6m. However, 27m_vs_30m has a larger number of units that need to cooperate.
C.3. MMM2
Figure 13. MMM2 replay.
The MMM2 is a scenario with a diverse set of units containing 1 medivac (medic), 2 marauders, and 7 marines fighting against a team with 1 more marauder and 1 more marine. Medivac cannot attack, but can heal units as long as it has energy. MACTAS ground units focus fire on enemy units and eliminate them one by one, and fall back when their health is low. The enemy medivac does not change healed unit, while the MACTAS’ medivac changes healed unit to the one with a critically low health or to the currently attacked unit. Units trained with MACTAS move medivac to the front line and focus enemy fire on it, and when its health level is low, retreat it. MACTAS’ medivac taking enemy fire secures outnumbered marines and marauders. Typically, human players try to save their medivacs because they are used for healing units between fights during a regular game. Only under certain conditions do they sacrifice their medivacs; for example, when they are breaking through an opponent’s line of defense to focus enemy fire on the medivacs and away from damage-dealing units. In this scenario, we can observe similar behavior in the MACTAS medivac control, which results in winning combat against an opponent who has a numerical advantage.
C.4. 6h_vs_8z
Figure 14. 6h_vs_8z replay.